# GraphCheck: Multi-Path Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck

### v2

@@ Abstract
Automated fact-checking aims to assess the truthfulness of textual claims based on relevant evidence. However, verifying complex claims that require multi-hop reasoning remains a significant challenge. We propose GraphCheck, a novel framework that transforms claims into entity-relationship graphs for structured and systematic verification. By explicitly modeling both explicit and latent entities and exploring multiple reasoning paths, GraphCheck improves verification robustness. While GraphCheck excels in complex scenarios, it may be unnecessarily elaborate for simpler claims. To address this, we introduce DP-GraphCheck, a variant that employs a lightweight strategy selector to adaptively choose between direct prompting and GraphCheck. This selective mechanism improves both accuracy and efficiency by applying the appropriate level of reasoning to each claim. Experiments on the HOVER and EX-FEVER datasets demonstrate that our approach outperforms existing methods, particularly on multi-hop claims. Moreover, the strategy selection mechanism in DP-GraphCheck generalizes well to other fact-checking pipelines, highlighting the versatility of our framework.
[…]
tcb@breakable
GraphCheck: Multi-Path Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck
@@ 1 Introduction
Automated fact-checking is a task that assesses the truthfulness of claims based on relevant evidence. With a standard pipeline that includes claim detection, evidence retrieval, and veracity assessment, automated systems enhance efficiency and accuracy in fact-checking Guo et al. (2022). However, verifying complex claims that require multi-hop reasoning remains a significant challenge. Such claims often consist of interwoven subclaims, making them difficult to verify at once. Also, relevant evidence is likely dispersed across multiple documents, complicating the retrieval process (Jiang et al., 2020; Aly et al., 2021). Another key obstacle is latent entities—references not explicitly stated in the text.
@@
For example, in the claim “The musician, who is part of Tall Birds, is a percussionist for a band that formed in Issaquah, Washington”, the phrases “The musician” and “a band” are latent entities. While these phrases correspond to specific entities (Davey Brozowski and Modest Mouse, respectively), the claim only implies relationships to these entities without explicitly revealing them. Identifying latent entities is crucial, as they can provide pivotal information for evidence retrieval and claim verification. Also, the order of latent entity identification matters, as some entities can be inferred more easily due to stronger contextual clues, which in turn help identify others as well. Conversely, initially misidentifying a challenging entity introduces false information, hindering subsequent identification steps. Thus, latent entity identification plays an important role in fact-checking.
Recent work has examined the application of large language models (LLMs) for verifying multi-hop claims without task-specific training, relying solely on few-shot in-context prompting (Brown et al., 2020; Dmonte et al., 2024). While this setting is attractive for its scalability and generality, it introduces unique challenges. Prior approaches typically leverage LLMs to generate verification paths by identifying “check-worthy” components Guo et al. (2022), such as subclaims and question-answering steps for latent entity identification (Pan et al., 2023b, a; Wang and Shu, 2023). However, due to the inherent ambiguity of check-worthiness, verification paths often lack granularity or omit key components. Moreover, these approaches typically rely on an LLM-generated, fixed verification path that may not be optimal, ultimately limiting verification accuracy.
@@
Figure 1: Overview of GraphCheck. Compared to prior approaches, GraphCheck offers a comprehensive yet fine-grained claim decomposition. It also systematically explores multiple paths instead of following a fixed path. The overall process consists of three steps: (1) A claim is converted into a structured entity-relation graph in which both explicit and latent entities are represented. (2) Latent entities are then identified through text infilling. Multiple infilling paths are explored via graph traversal, resulting in multiple infilled graphs. (3) For each infilled graph, every triplet is individually verified. The claim is classified as Supported if at least one path yields a graph in which all triplets are verified as Supported; otherwise, the claim is classified as Not Supported.
@@
While GraphCheck thoroughly verifies complex claims, it may be unnecessarily elaborate for simpler cases. Some relatively simple claims can be verified more efficiently and effectively through direct prompting (DP or Direct), where the LLM directly assesses a claim’s truthfulness based on retrieved documents. To leverage the complementary strengths of Direct and GraphCheck, we introduce DP-GraphCheck, which employs a lightweight strategy selector to adaptively choose between the two methods. Notably, this strategy selector is modular and can be easily integrated into diverse fact-checking systems beyond GraphCheck.
Experimental results on multi-hop fact-checking datasets (Jiang et al., 2020; Aly et al., 2021) demonstrate that both GraphCheck and DP-GraphCheck outperform existing methods on complex multi-hop verification tasks, highlighting the effectiveness of structuring claims as entity-relationship graphs for comprehensive and systematic fact-checking. Furthermore, we find that the strategy selector improves performance not only in GraphCheck but also when applied to other baseline fact-checking systems, underscoring its broad applicability.
@@ 2 Related Work
LLM-based Fact-Checking
Traditional fact-checking has relied on rule-based systems or supervised models trained on annotated datasets. With the rise of large language models (LLMs), the field has shifted towards more flexible, context-aware reasoning (Guo et al., 2022; Dmonte et al., 2024). LLM-based methods utilize retrieval-augmented generation (RAG) (Lewis et al., 2021; Gao et al., 2024) and in-context learning Brown et al. (2020) for veracity prediction.
For multi-hop fact-checking, where reasoning across multiple evidence pieces is required, LLM-based strategies include iterative RAG Shao et al. (2023) and claim decomposition (Chen and Shu, 2023; Zhang and Gao, 2023; Pan et al., 2023b). GraphCheck introduces a novel approach, leveraging a graph structure for claim decomposition to enable comprehensive multi-hop verification.
@@
Identifying latent entities is key to improving evidence retrieval and verification accuracy. While Prior work has used question-answering (Pan et al., 2023b, a; Wang and Shu, 2023) or RAG Shao et al. (2023) to tackle this, GraphCheck constructs an entity-relationship graph where latent entities are explicitly represented as nodes and identified through text infilling Zhu et al. (2019).
Existing methods also rely on model-driven ordering for latent entity identification.  Shao et al. (2023) and  Pan et al. (2023a) adopt an iterative approach where prior outputs guide subsequent steps, forming a linear sequence.  Pan et al. (2023b) generates an entire reasoning sequence in a single LLM API call and aggregates results from multiple paths generated by additional calls. In contrast, GraphCheck systematically explores multiple identification paths within a graph generated by a single API call, ensuring efficient and robust verification.
@@
Graph-based models have been utilized in fact-checking to capture relationships between entities in claims and evidence.  Zhong et al. (2020) employed semantic role labeling (SRL) to construct graphs from evidence and applied graph convolutional networks (GCNs) and graph attention networks (GATs) for logical inference.  Yuan and Vlachos (2024) extracted triplets using an OpenIE model and verified them through a natural language inference (NLI) model. Unlike prior approaches, GraphCheck constructs entity-relationship graphs that explicitly represent latent entities as placeholder nodes, leveraging the in-context learning capabilities of LLMs.
@@ 3.1 GraphCheck: Graph Construction
GraphCheck first transforms a claim into an entity-relationship graph. Latent entities within the claim are detected and represented as placeholders, forming distinct nodes. To construct the graph, we leverage an LLM guided by instructions that include predefined rules and few-shot examples. The key instructions are as follows:
  • Decompose the claim into fact triplets, each serving as a basic unit of information.
  • Format a triplet as ‘subject [SEP] relation [SEP] object’, where the subject and object are noun phrases, and the relation is verb phrase, forming a complete sentence.
  • Detect and represent latent entities using placeholders, e.g., (ENT1), (ENT2).
As shown in the part of Figure 1 that explains the “(1) Graph Construction” step of GraphCheck, the generated graph consists of two sections:
@@
While triplets in both section serve as subclaims requiring verification, the # Latent Entities section is set apart to ensure that contextual meaning is preserved when placeholders are introduced. For instance, in the example shown in Figure 1, replacing “the musician” with a placeholder (ENT1) could result in a loss of information. To prevent this, placeholders are explicitly mapped to their corresponding references in the seperated section.
@@ 3.2 GraphCheck: Latent Entity Infilling
Once the entity-relationship graph is constructed, latent entities are identified sequentially through text infilling. The process begins by retrieving top-$k$ documents relevant to the target latent entity. To ensure sufficient context, a retrieval query is generated by concatenating all triplets that include the target entity and exclude any other unidentified latent entities. The retrieved documents, along with the same set of triplets, are then used as input to the text infilling model. Within this input, the target latent entity is represented by a special token. A more detailed description of the infilling process is provided in Appendix B.2.
@@
Multiple Path Exploration
For claims containing multiple latent entities, various identification paths are explored. In the example shown in Figure 1, two possible paths—(ENT1$\rightarrow$(ENT2), (ENT2$\rightarrow$(ENT1)—are considered. To manage computational complexity, up to $\bar{P}$ paths are randomly sampled if the total number of possible sequences exceeds this limit.
Exploring various identification paths is crucial, as some paths may be more effective than others. Certain latent entities can be identified more easily due to stronger contextual clues, and identifying them first can provide additional context for subsequent identifications. For example, in Figure 1, identifying (ENT2) is easier than identifying (ENT1) because “Issaquah, Washington” provides a more salient retrieval cue than “Tall Birds”, which may introduce ambiguity between a rock band and actual birds. As a result, when (ENT2) is identified first in Path 2, both entities are correctly infilled, whereas Path 1 fails to infill (ENT1) correctly.
However, automatically finding the optimal identification path is a challenging task. To address this, our method systematically explores multiple paths rather than relying on model-driven path planning, increasing the likelihood of finding the most effective path.
@@ 3.3 GraphCheck: Fact Triplet Verification
After infilling, each triplet in the graph is independently verified. For each infilled triplet $t^{\prime}$, the top-$k$ documents are retrieved from the corpus using $t^{\prime}$ as the retrieval query. The verifier assesses the veracity of $t^{\prime}$ using two types of evidence inputs: (i) the concatenation of the top-$k$ documents, and (ii) each of the top-$k$ documents individually, resulting in a total of $k+1$ inputs. If any of these inputs yields a Supported judgment, $t^{\prime}$ is classified as Supported; otherwise, it is classified as Not Supported.
@@
A latent entity identification path produces a fully infilled graph consisting of triplets $\{t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}\}$. The path is classified as Supported if all triplets are classified as Supported; otherwise, it is classified as Not Supported.
@@
Since multiple identification paths can be explored, GraphCheck performs triplet-level verification for each path independently. A claim is ultimately classified as Supported if at least one path is classified as Supported; otherwise, it is classified as Not Supported.
This classification approach accounts for potential errors in latent entity infilling. A claim classified as Not Supported in a single identification path does not necessarily indicate that the claim itself is false; rather, it may result from improperly identified latent entities. Exploring multiple identification paths increases the likelihood of correct verification, as a claim is successfully verified if at least one path accurately identifies latent entities, thereby supporting robust fact-checking.

### v3

# GraphCheck: Multipath Fact-Checking with Entity-Relationship GraphsCode available at: https://github.com/windowh1/graphcheck

@@ Abstract
Automated fact-checking aims to assess the truthfulness of textual claims based on relevant evidence. However, verifying complex claims that require multi-hop reasoning remains a significant challenge. We propose GraphCheck, a novel framework that transforms claims into entity-relationship graphs for structured and systematic fact-checking. By explicitly modeling both explicit and latent entities and exploring multiple reasoning paths, GraphCheck enhances verification robustness. While GraphCheck excels in complex scenarios, it may be unnecessarily elaborate for simpler claims. To address this, we introduce DP-GraphCheck, a variant that employs a lightweight strategy selector to choose between direct prompting and GraphCheck adaptively. This selective mechanism improves both accuracy and efficiency by applying the appropriate level of reasoning to each claim. Experiments on the HOVER and EX-FEVER datasets demonstrate that our approach outperforms existing methods in verification accuracy, while achieving strong computational efficiency despite its multipath exploration. Moreover, the strategy selection mechanism in DP-GraphCheck generalizes well to other fact-checking pipelines, highlighting the broad applicability of our framework.
GraphCheck: Multipath Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck
@@ 1 Introduction
Automated fact-checking is a task that assesses the truthfulness of claims based on relevant evidence. With a standard pipeline that includes claim detection, evidence retrieval, and veracity assessment, automated systems enhance efficiency and accuracy in fact-checking Guo et al. (2022). However, verifying complex claims that require multi-hop reasoning remains a significant challenge. Such claims often consist of interwoven subclaims, making them difficult to verify at once. Also, relevant evidence is likely dispersed across multiple documents, complicating the retrieval process (Jiang et al., 2020; Ma et al., 2024).
@@
Another key obstacle is the existence of latent entities—references not explicitly stated in the text. For example, in the claim “The musician, who is part of Tall Birds, is a percussionist for a band that formed in Issaquah, Washington”, the phrases “The musician” and “a band” are latent entities. While these phrases correspond to specific entities (Davey Brozowski and Modest Mouse, respectively), the claim only implies relationships to these entities without explicitly revealing them. Identifying latent entities is crucial, as they can provide pivotal information for evidence retrieval and claim verification. Also, the order of latent entity identification matters, as some entities can be inferred more easily due to stronger contextual clues, which in turn help identify others as well. Conversely, initially misidentifying a challenging entity introduces false information, hindering subsequent identification steps. Thus, latent entity identification plays an important role in fact-checking.
@@
Figure 1: Overview of GraphCheck. Compared to prior approaches, GraphCheck offers a comprehensive yet fine-grained claim decomposition. It also systematically explores multiple paths instead of following a fixed path. The overall process consists of three steps: (1) A claim is converted into a structured entity-relationship graph in which both explicit and latent entities are represented. (2) Latent entities are then identified through text infilling. Multiple infilling paths are explored via graph traversal, resulting in multiple infilled graphs. (3) For each infilled graph, every triplet is individually verified. The claim is predicted as Supported if at least one path yields a graph in which all triplets are verified as Supported; otherwise, the claim is predicted as Not Supported.
Recent work has examined the application of large language models (LLMs) for verifying multi-hop claims, relying solely on few-shot prompting without additional task-specific training (Brown et al., 2020; Dmonte et al., 2024). This approach is appealing due to its scalability and generality, yet it still faces notable limitations. Prior methods leverage LLMs to generate verification paths by identifying “check-worthy” components Guo et al. (2022), such as subclaims to be verified and question-answering steps for latent entity identification (Pan et al., 2023b, a; Wang and Shu, 2023). However, due to the inherent ambiguity of check-worthiness, verification paths often lack granularity or omit key components. Moreover, these methods typically rely on an LLM-generated, fixed verification path that may not be optimal, ultimately limiting verification accuracy.
@@
While GraphCheck thoroughly verifies complex claims, it may be unnecessarily elaborate for simpler cases. Some relatively simple claims can be verified more efficiently and effectively through direct prompting (DP or Direct), where the LLM directly assesses a claim’s truthfulness based on relevant documents. To leverage the complementary strengths of Direct and GraphCheck, we introduce DP-GraphCheck, which employs a lightweight strategy selector to choose between the two methods adaptively. Notably, this strategy selector is modular and can be easily integrated into diverse fact-checking systems beyond GraphCheck.
Experimental results on multi-hop fact-checking datasets (Jiang et al., 2020; Ma et al., 2024) demonstrate that both GraphCheck and DP-GraphCheck outperform existing methods on the complex multi-hop fact-checking task, highlighting the effectiveness of structuring claims as entity-relationship graphs. Despite its multipath exploration, GraphCheck achieves superior computational efficiency compared to multipath baselines. Moreover, the strategy selector improves performance not only within GraphCheck but also across other baseline fact-checking systems, underscoring its broad applicability.
@@ 2 Related Work
Multi-hop Fact-Checking
Fact-checking involves assessing the veracity of claims based on supporting evidence. Early research primarily focuses on single-hop fact-checking, where the evidence necessary to validate a claim is contained within a single document or passage (Vlachos and Riedel, 2014; Wang, 2017; Thorne et al., 2018). To better reflect real-world situations, where claim verification often depends on dispersed or interconnected information, subsequent work proposes multi-hop fact-checking, which demands reasoning across multiple pieces of evidence (Jiang et al., 2020; Ma et al., 2024; Aly et al., 2021).
Early approaches to multi-hop fact-checking rely on task-specific supervised training of neural models with annotated datasets (Jiang et al., 2020; Ostrowski et al., 2021; Khattab et al., 2022). While these methods demonstrate strong in-domain performance and maintain computational efficiency during inference through lightweight architectures, they are fundamentally limited by the high cost of large-scale annotation and poor cross-domain generalization.
The advent of large language models (LLMs) has shifted the field toward more flexible, context-aware reasoning (Guo et al., 2022; Dmonte et al., 2024). LLM-based approaches often combine retrieval-augmented generation (RAG) (Lewis et al., 2021; Gao et al., 2024) with in-context learning Brown et al. (2020) to support veracity prediction without additional task-specific training. Recent work has explored various strategies, such as iterative RAG Shao et al. (2023) and claim decomposition (Chen and Shu, 2023; Zhang and Gao, 2023; Pan et al., 2023b). Building on these advances, GraphCheck introduces a graph-based decomposition approach that structures claims into entities and relations, enabling more comprehensive multi-hop fact-checking.
@@
Identifying latent entities is crucial for improving evidence retrieval and verification accuracy in multi-hop fact-checking. Previous approaches have addressed this challenge using question-answering frameworks (Pan et al., 2023a, b; Wang and Shu, 2023). GraphCheck takes a different approach by constructing an entity-relationship graph where latent entities are represented as placeholder nodes and identified through text infilling Zhu et al. (2019).
Regarding the ordering of latent entity identification, existing methods follow model-driven approaches with distinct strategies. Specifically,  Shao et al. (2023) and  Pan et al. (2023a) employ iterative processes where each step builds on previous outputs, creating linear reasoning chains, while  Pan et al. (2023b) generates a complete reasoning path within a single LLM call and aggregates results across multiple generated paths. In contrast, GraphCheck systematically explores multiple identification paths within a graph structure constructed through a single API call, thereby achieving both efficiency and robustness.
@@
Recent advances in fact-checking have increasingly adopted graph-based methods to represent structured relationships in textual claims or evidence. GEAR Zhou et al. (2019) models each evidence sentence as a node and applies graph neural networks to aggregate information across sentences. KGAT Liu et al. (2020) represents claim–evidence pairs as nodes and applies kernel-based attention over the graph.
Other approaches focus on entity-level relationships. Zhong et al. Zhong et al. (2020) employ semantic role labeling to construct graphs from evidence, while Yuan and Vlachos Yuan and Vlachos (2024) extract triplets from claims with OpenIE models and verify them using NLI models. GraphCheck takes a distinct approach by constructing entity–relationship graphs that represent latent entities as placeholder nodes, leveraging the in-context learning capabilities of LLMs.
@@ 3.1 GraphCheck: Graph Construction
GraphCheck first transforms a claim into an entity-relationship graph. Latent entities within the claim are detected and represented as placeholder nodes. To construct the graph, we leverage an LLM guided by instructions that include predefined rules and few-shot examples (Appendix H.1). The key instructions are as follows:
  • Detect and represent latent entities using placeholders (e.g., (ENT1), (ENT2)).
  • Decompose the claim into fact triplets (e.g., subject [SEP] relation [SEP] object), each serving as a basic unit of information.
As shown in the part of Figure 1 that explains the (1) Graph Construction step of GraphCheck, the generated graph consists of two sections:
@@
While triplets in both sections serve as subclaims requiring verification, the # Latent Entities section is set apart to ensure that contextual meaning is preserved when placeholders are introduced. For instance, in the example shown in Figure 1, replacing “the musician” with a placeholder (ENT1) could result in a loss of information. To prevent this, placeholders are explicitly mapped to their corresponding references in the separated section.
@@ 3.2 GraphCheck: Latent Entity Infilling
Once the entity-relationship graph is constructed, latent entities are identified sequentially through text infilling. The process begins by retrieving top-$k$ documents relevant to the target latent entity. To ensure sufficient context, a retrieval query is formulated by concatenating all triplets that include the target entity and exclude any other unidentified latent entities. The retrieved documents, along with the same set of triplets, are then used as input to the text infilling model. A more detailed description of the infilling process is provided in Appendix H.2.
@@
Multipath Exploration
[…]
Exploring various identification paths is important because some paths can be more effective than others. Some latent entities are identified more easily due to stronger contextual clues, and identifying them first can provide additional context for subsequent identifications. For example, in Figure 1, identifying (ENT2) is easier than identifying (ENT1) because “Issaquah, Washington” provides a more salient retrieval cue than “Tall Birds”, which may introduce ambiguity between a rock band and actual birds. As a result, when (ENT2) is identified first in Path 2, both entities are correctly infilled, whereas Path 1 fails to infill (ENT1) correctly.
However, automatically finding the optimal path is challenging. To address this, our method systematically explores multiple paths instead of relying on model-driven planning, thereby increasing the likelihood of finding the most effective path.
@@ 3.3 GraphCheck: Fact Triplet Verification
After infilling, each triplet in the graph is independently verified. For each infilled triplet $t^{\prime}$, the top-$k$ documents are retrieved from the corpus using $t^{\prime}$ as the retrieval query. The verifier assesses the veracity of $t^{\prime}$ using k+1k{+}1 evidence inputs: (i) the concatenation of the top-$k$ documents, and (ii) each of the top-$k$ documents individually. If any of these inputs yields a Supported judgment, $t^{\prime}$ is classified as Supported; otherwise, it is classified as Not Supported.
@@
A latent entity identification path produces a fully infilled graph consisting of triplets $\{t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}\}$. The path is classified as Supported if all triplets in the graph are classified as Supported; otherwise, it is classified as Not Supported.
@@
Since multiple identification paths can be explored, GraphCheck performs triplet-level verification for each path independently. A claim is ultimately classified as Supported if at least one path is classified as Supported; otherwise, it is ultimately classified as Not Supported.
This classification approach accounts for potential errors in latent entity infilling. A claim classified as Not Supported in a single path does not necessarily indicate that the claim itself is false; rather, it may stem from incorrectly identified latent entities. By exploring multiple identification paths, GraphCheck increases the likelihood of accurate verification, as the claim veracity can be reliably assessed if at least one path correctly identifies latent entities.